OpenStudy (anonymous):

f(x)=IxI+5 Find the Range. Write in interval notation

OpenStudy (anonymous):

Range f(x)>=5

OpenStudy (anonymous):

how is that written in interval notation?

OpenStudy (anonymous):

[5,infinity)

OpenStudy (anonymous):

i mean use the symbol for infinity

OpenStudy (anonymous):

so would IxI-6 be [-6,oo)?

OpenStudy (anonymous):

\[|x| \ge 0\]\[|x| - 6 \geq -6\]in interval notation, \[[-6 , \infty)\]